OpenStudy (inspectorjoe):

Im trying to solve this problem and my answer keep coming up in decimals. need help 2^6X2^-3/20^10/2-8

OpenStudy (inspectorjoe):

\[2^{6}\times2^{-3}\div2^{10}\div2^{-8}\]

Nnesha (nnesha):

don't use calculator. you will not get decimal answer.

OpenStudy (inspectorjoe):

I'm not which means I must be doing something wrong

Nnesha (nnesha):

ohh nvm then show your work please don't even touch the base just move around the exponents

Nnesha (nnesha):

exponent rule you can't have the negative exponent \[x^{-m}=\frac{ 1 }{ x^m }\] flip the fraction to make it positive exponent

Nnesha (nnesha):

\[\huge\rm \frac{ x^{-m} }{ 1 }=\frac{ 1 }{ x^m }\]

OpenStudy (inspectorjoe):

are we talking about an imaginary base under the entire problem ? sorry

Nnesha (nnesha):

imaginary base ??? :o

Nnesha (nnesha):

\[\huge\rm \frac{ 2^6 \times 2^{-3} }{ \frac{ 2^{10} }{ 2^{-8}} }\] this is your question right where `2` is the base

OpenStudy (inspectorjoe):

NO

OpenStudy (inspectorjoe):

refer to the initial post showing the actual problem

Nnesha (nnesha):

\[\huge\rm \frac{\color{reD}{ 2^6 \times 2^{-3}} }{ \frac{ 2^{10} }{ 2^{-8}} }\] solve the red part first when you `multiply` same bases you should `add` their exponents

OpenStudy (inspectorjoe):

okay what about the division ? my results now come back to 1^1

Nnesha (nnesha):

alright \[\rm \color{Red}{2^6 \times 2^{-3} }\div 2^{10} \div 2^{-8}\]

OpenStudy (inspectorjoe):

Yes

Nnesha (nnesha):

same thing. :=) solve the red part first

OpenStudy (inspectorjoe):

okay than divide from left to right subtracting the exponents right?

OpenStudy (inspectorjoe):

okay so my new answer is 4^{27}

OpenStudy (inspectorjoe):

4^27

Nnesha (nnesha):

how did you get 4 ?

OpenStudy (zzr0ck3r):

there is nothing wrong with decimals

OpenStudy (inspectorjoe):

2X2

Nnesha (nnesha):

just play with the exponents don't even touch the base

Nnesha (nnesha):

multiply the same u would add their exponents alright \[\rm \color{Red}{2^{6+(-3)} }\div 2^{10} \div 2^{-8}\] like this

OpenStudy (inspectorjoe):

So leave the base and add and subtract the exponents? new answer\[2^{-36}\]

OpenStudy (inspectorjoe):

correction sorry wrong numbers

Nnesha (nnesha):

no.. i said when you `multiply` the same bases then u should add \[\huge\rm \color{Red}{2^6 \color{blue}{\times }2^{-3} }\div 2^{10} \div 2^{-8}\] here are multiplying 2^6 times 2^{-3 that's why we should add the base

OpenStudy (inspectorjoe):

2^21 final answer

Nnesha (nnesha):

\[\huge\rm \color{black}{2^{6+(-3)} }\color{red}{\div} 2^{10} \color{red}{\div }2^{-8}\] different rule when you divide same bases \[\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (inspectorjoe):

okay so my results should be 2^-15

Nnesha (nnesha):

how did you get -15 ?

OpenStudy (inspectorjoe):

okay I first did 6+-3= 3 than I did 10- -8= 18 3-18=-15?

Nnesha (nnesha):

thanks for showing your work much appreciated ! \[\huge\rm \frac{ \color{ReD}{\frac{ 2^3 }{ 2^{10}}} }{ 2^{-8} }\] first solve the red part we should move the 10 to the top first and then

Nnesha (nnesha):

\[\huge\rm \frac{ 2^{3+(-10)} }{ 2^{-8}}\] try to solve now!

OpenStudy (inspectorjoe):

2^1? final answer, I did 3+-10= -7 than -7 - -8= 1 2^1

Nnesha (nnesha):

YAYAY!!

OpenStudy (inspectorjoe):

Really I though I got it wrong waohh

OpenStudy (inspectorjoe):

but how did we went from the original problem to this?

OpenStudy (inspectorjoe):

final one?

Nnesha (nnesha):

we used two exponents rule! \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\] divide same base \[\huge\rm x^m \times x^n = x^{m \times n}\] when you multiply same bsses

OpenStudy (inspectorjoe):

holly cow for some reason I feel lost a bit. questions can you give a practice problem to solve for you.

Nnesha (nnesha):

sure sure \[\huge\rm \frac{\frac{ 3^6 \times 3^7 }{ 3^4} }{ 3^{-5} }\]

OpenStudy (inspectorjoe):

give me a sec to resolve thanks

OpenStudy (zzr0ck3r):

I want to make one thing clear: you are of course allowed to have negative exponents it is just that your teacher wants to make sure you can go back and fourth.

Nnesha (nnesha):

so why we can't leave the negative exponent when it says `simplify`

Nnesha (nnesha):

:=)

Nnesha (nnesha):

actually sometimes we could , depends on the answer choices :P haha

OpenStudy (anonymous):

if you need extra help you can use this site https://mathway.com/

OpenStudy (inspectorjoe):

thanks pinkiepug I will

OpenStudy (anonymous):

your welcome ^.^

OpenStudy (inspectorjoe):

okay here is my answer and problem